This is a word problem for inverse variations.
The cost per person to rent a mountain cabin is inversely proportional to the number of people who share the rent. If the cost is $36 per person when 5 people share, what is the cost per person when 8 people share?
using the formula y=k/x

Answer by stanbon(75887)

The cost per person to rent a mountain cabin is inversely proportional to the number of people who share the rent.
This means y=x/k
You need to find "k".
------------
If the cost is $36 per person when 5 people share
So 36=k/5
k=180
Now we know the formula is y=36/5
---------------
what is the cost per person when 8 people share?
Use the formula to find the cost:
cost=180/8
cost=$22.50 per person
Cheers,
stan H.